OpenStudy (anonymous):

At summer camp, 20 campers try horseback riding or climbing. If 13 campers try horseback and 15 campers try climbing, how many campers try both activities?

OpenStudy (anonymous):

@johnweldon1993 @kelliegirl33 @Jonathan1997 @RadEn @petiteme @agent0smith anybody ?

OpenStudy (anonymous):

u should use venn diagram

OpenStudy (anonymous):

I was gonna say that

OpenStudy (agent0smith):

 20 = 15-x + 13-x + x solve for x

OpenStudy (raden):

alternative : n(A⋃B) = n(A) + n(B) - n(a⋂b) 20 = 13 + 15 - n(a⋂b) solve for n(a⋂b) ?

OpenStudy (anonymous):

woukld it be 6 ?

OpenStudy (anonymous):

or 8?

OpenStudy (agent0smith):

Which one? Are you sure?

OpenStudy (anonymous):

no im not sure